OpenStudy (devinmartin0923):

Need Algebra 2 Help!

OpenStudy (devinmartin0923):

Use the four functions below for this question:

OpenStudy (devinmartin0923):

Compare and contrast these four functions using complete sentences. Focus on the slope and y-intercept first, and then on any additional properties of each function.

OpenStudy (mathmate):

hints: In a function of the form y=ax+b b is the y-intercept (because y=b when x=0) and a is the slope. The slope determines if the y is increasing or decreasing as x increases. a>0 means slope is positive (increasing) a<0 means slope is negative (decreasing) a=0 means the line is horizontal.

OpenStudy (devinmartin0923):

All slopes are negative correct?

OpenStudy (devinmartin0923):

Confused what its asking for f(x) @mathmate

OpenStudy (devinmartin0923):

@legomyego180

OpenStudy (flvskidd):

idk sry

OpenStudy (mathmate):

y and f(x) are kind of synonyms. y is used in high-school, and f(x) is used thereafter. f(x) specifies what the independent variable is, and for y, the variable is understood to be x. We could write equally y=ax+b or f(x)=ax+b where f(x) is actually the y-coordinate on the graph.

OpenStudy (devinmartin0923):

this is f(x) = Jesse got $5 from his dad and he makes $2 each time he walks the neighbor's dog.

OpenStudy (mathmate):

Not all four functions have negative slopes. When f(x) (or y) increases when x increases, the slope is positive. For example, f(x)=2x-3 the slope is +2, so the slope is positive. f(x)=-2x-3 the slope is -2, so the slope is negative. "b" has no incidence on the value of the slope. It just shifts the line up or down. Please reread the significance of a and b in f(x)=ax+b.

OpenStudy (devinmartin0923):

The first two are positive and the last two are negative slopes?

OpenStudy (devinmartin0923):

or is the first and last positive and the middle negative, this is so confusing to me

OpenStudy (devinmartin0923):

@mathmate

OpenStudy (mathmate):

Remember, a positive slope means that the function f(x) (or g(x), h(x), i(x)...) \(increases\) as x increases. For example, x 1 3 7 y 3 7 15 is an increasing function because as x goes from 1 to 3 (increase), y goes from 3 to 7 (increase).

OpenStudy (mathmate):

For j(x) Did you graph it on a calculator, or did you make a table? If you made a table, can you post it?

OpenStudy (devinmartin0923):

everything is increasing except for h?

OpenStudy (devinmartin0923):

i mad an error i see the graph is increasing

OpenStudy (mathmate):

You're giving me four different answers in 5 minutes. Tell me how you work, on a computer, on paper or with a calculator? Work with one function at a time, make sure you're correct before trying others. This way, you confirm your concept, and not confuse yourself with different combination of answers.

OpenStudy (devinmartin0923):

I dont know how to do the first two

OpenStudy (mathmate):

Make a table, and if the computer cannot handle it, use paper. Using paper has the advantage that the next time you see a table, you can visualize what it \(should\) look like. With the computer or calculator, once you press reset, nothing is left in your mind, or, you hardly learned anything.

OpenStudy (devinmartin0923):

2x + 5 is the first one

OpenStudy (devinmartin0923):

the first and last are increasing slopes and the middle two are negative. with your help this is what ive concluded.

OpenStudy (mathmate):

I agree with your conclusions, well done! Now, you may want to proceed to the y-intercepts.

OpenStudy (mathmate):

Also, you may want to find the slope of each, not just the sign.

OpenStudy (devinmartin0923):

The first and last y intercept is 5?

OpenStudy (devinmartin0923):

the first and last are the same but with swapped - +

OpenStudy (mathmate):

So slope / y-intercept f + 5 g - h - j + -5 Can you complete the rest?